# Matching Markets Meet LLMs: Algorithmic Reasoning with Ranked Preferences

### v1

@@ Abstract
The rise of Large Language Models (LLMs) has driven progress in reasoning tasks—from program synthesis to scientific hypothesis generation—yet their ability to handle ranked preferences and structured algorithms in combinatorial domains remains underexplored. We study matching markets, a core framework behind applications like resource allocation and ride-sharing, which require reconciling individual ranked preferences to ensure stable outcomes. We evaluate seven state-of-the-art models on a hierarchy of preference-based reasoning tasks—ranging from stable-matching generation to instability detection, instability resolution, and fine-grained preference queries—to systematically expose their logical and algorithmic limitations in handling ranked inputs. Surprisingly, even top-performing models with advanced reasoning struggle to resolve instability in large markets, often failing to identify blocking pairs or execute algorithms iteratively. We further show that parameter-efficient fine-tuning (LoRA) significantly improves performance in small markets, but fails to bring about a similar improvement on large instances, suggesting the need for more sophisticated strategies to improve LLMs’ reasoning with larger-context inputs.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has positioned them as integral components in a wide range of reasoning-intensive tasks such as program synthesis, logical inference, mathematical problem solving, and scientific hypothesis generation, highlighting the importance of structured problem-solving capabilities. Despite their recent success in symbolic and logical reasoning, their capacity to reason over ranked preferences and to execute structured algorithms within combinatorial domains remains largely unexplored. Preference reasoning constitutes a foundational component in numerous domains, including economic contexts—e.g., auctions, voting systems, and market design—and in the architecture of pre-trained generative models using Reinforcement Learning from Human Feedback (RLHF) to capture and internalize human value judgments. These methods often have to execute algorithms on a large number of preference lists (either pairwise, partial, or complete rankings) to aggregate the rankings through constitutional AI [2] or social choice theory [15].
Despite substantial progress, reasoning over preferences remains a non-trivial endeavor: ensuring transitivity [73, 62], accurately augmenting ordinal rankings [23], and achieving coherent value alignment pose significant challenges. Without robust mechanisms for preference elicitation and the capacity to execute the requisite combinatorial procedures, even state-of-the-art LLMs may produce outputs that diverge from true human preferences [33] or fail to satisfy desirable properties [24].
We consider matching markets, a domain that constitutes a fundamental class of problems underlying diverse applications—from healthcare resource allocation to ride-sharing platforms and recommender systems—and demand accurate comprehension of individual preferences and reconciliation of conflicting choices to guarantee system-wide stability. Matching markets are a compelling domain for studying reasoning in AI models for two key reasons: First, they provide a structured platform for evaluating reasoning over ranked preferences and algorithmic thinking. Second, LLMs are increasingly utilized as black-box systems in a variety of economic, social, or medical settings to inform automated screening in recruitment pipelines [28], investigating market behavior [38], market clearing in ride-hailing platforms [41], and in general simulating economic interactions [32].
@@ 1.1 Our Results
We focus on four preference-based tasks: (i) generating stable solutions, requiring LLMs to produce stable matchings directly from ranked inputs, (ii) instability resolution, demanding preference reasoning to transform unstable matchings to stable ones, (iii) instability detection, in which models detect blocking pairs within a proposed solution, and (iv) preference reasoning, assessing nuanced query answering over ranked lists. We evaluate seven large language models with varying reasoning capabilities, including basic models, those with some limited reasoning, and advanced reasoning models. The following points highlight our contributions and core findings:
Benchmark. We introduce a benchmark with instances and questions aimed at evaluating the above tasks involving reasoning over ranked preferences. These tasks are categorized into three levels of difficulty—Easy, Medium, and Hard—based on problem size. Each task utilizes ranked preferences sampled from two statistical distribution models: Impartial Culture (IC) and Master List (ML).
@@
Instability Detections and Resolution. We find that LLMs frequently make mistakes in determining whether solutions are stable, with hallucinations about blocking pairs being the most common among basic models. Additionally, LLMs’ ability to correct unstable solutions is (at best) as good as their ability to generate them from scratch, in some cases making the provided incorrect solutions worse.
@@ 1.2 Related Work
Reasoning Capabilities of LLMs. Mathematical problem solving has been a key area of focus in evaluating the reasoning ability of LLMs, through a variety of benchmarks such as [31, 30, 14, 56]. LLMs have also demonstrated remarkable capabilities on coding benchmarks such as SWE-Bench [40] and CodeForces. As SOTA benchmark scores improve, recent work studies whether these improvements reflect genuine logical reasoning through benchmarks assessing logical consistency [49] and rule understanding/execution/planning [29]. Furthermore, the recent rise of LLM agents has increased interest in benchmarking LLMs’ causal reasoning [12] and strategic planning abilities [20, 64, 39]. Additionally, the emergence of reasoning models has led to benchmarks evaluating these models’ improved reasoning and planning abilities [10, 48].
Enhancing Reasoning Capabilities of LLMs. Specialized prompting strategies like Chain-of-Thought (CoT) [70], Tree-of-Thought (ToT) [75], and Graph-of-Thought (GoT) [3] have performance abilities on a variety of reasoning benchmarks. Fine-tuning has also been demonstrated to improve CoT in model outputs [76], as well as economic rationality [11] and abstract reasoning [72]. Additionally, instruction-tuning has been shown to enhance reasoning in several works [50, 69, 8, 47]. More advanced techniques build upon CoT [66, 78], or utilize multi-agent architectures that leverage cooperative LLMs [77, 67]. More recently, reinforcement-learning (GRPO) has been used to improve model reasoning, the most popular example being the Deepseek-R1 reasoning model [17].
LLMs in Social and Economic Decision Making. While still an emerging area of research, multiple works have focused on the collective decision-making capabilities of LLMs. One particular area of interest is the use of LLMs in preference elicitation [63, 35]. Fish et al. [24] benchmark the ability of models to learn and strategize in unknown economic environments using deliberate exploration. Another notable avenue of work is the study of how well LLMs can represent humans in collective decision-making, an understudied component of LLM alignment [74, 33].
@@ 2.1 Problem Formulation
A two-sided matching market consists of two disjoint sets of agents (e.g., riders and drivers, freelancers and job requesters, and content creators and ads) denoted by $M$ and $W$, where $|M|=|W|=n$. The preference list of an agent $i$, denoted by $\succ_{i}$, is a ranked order list over the agents on the other side. A preference profile, $\succ$, denotes the collection of preferences of all agents. We write $w_{1}\succ_{m}w_{2}$ and $m_{1}\succ_{w}m_{2}$ to denote that $m$ prefers $w_{1}$ to $w_{2}$ and $w$ prefers $m_{1}$ to $m_{2}$ respectively. In this paper, we primarily consider the standard model, which assumes a complete and strict preference list (no ties) and aims at finding a one-to-one matching between the agents in two sets. [^1] This is the standard model considered by the seminal works of Gale and Shapley [27] and Knuth [44].
Matching and Stability. A matching is a function $\mu:M\cup W\rightarrow M\cup W$ such that $\mu(m)\in W$ for all $m\in M$, $\mu(w)\in M$ for all $w\in W$, and $\mu(m)=w$ if and only if $\mu(w)=m$. Given a matching $\mu$, a blocking pair with respect to the preference profile $\succ$ is a pair $(m,w)$ who prefer each other over their assigned partners in $\mu$, i.e., $w\succ_{m}\mu(m)$ and $m\succ_{w}\mu(w)$. A matching is said to be stable if it does not have any blocking pairs. Given an instance of the problem, the set of all possible stable solutions forms a distributive lattice and can be exponential in size [44].
In their seminal work, Gale and Shapley [27] proposed an iterative procedure—the deferred acceptance algorithm (DA)—that always guarantees to find a stable solution. It proceeds by a series of proposals and rejections. In the initial proposal phase, each of the unmatched agents on one side (aka proposers) proposes to their favorite agent from the other side (aka receivers) according to their preference list. In the subsequent rejection phase, each agent on the receiving side tentatively accepts its preferred proposal, rejecting the others. The algorithm terminates when no further proposals can be made. The details of this algorithm can be found in Appendix C. The underlined solution in Example 1 is simultaneously optimal for the proposing side and pessimal for the receiving side [55]. We refer to the former as the Optimal matching and the latter as the Pessimal matching.
@@
${\boldsymbol{m_{1}}}\colon$ $\underline{w_{4}}$ $w_{3}$ $w_{1}^{*,\dagger}$ $w_{2}$ $\boldsymbol{w_{1}}\colon$ $m_{2}$ ${m_{1}^{*,\dagger}}$ $\underline{m_{3}}$ $m_{4}$
${\boldsymbol{m_{2}}}\colon$ $\underline{w_{3}}$ ${w_{4}^{\dagger}}$ $w_{2}^{*}$ $w_{1}$ ${\boldsymbol{w_{2}}}\colon$ $m_{2}^{*}$ $m_{3}$ ${\underline{m_{4}^{\dagger}}}$ $m_{1}$
${\boldsymbol{m_{3}}}\colon$ $\underline{w_{1}}$ $w_{3}^{\dagger}$ $w_{2}$ $w_{4}^{*}$ $\boldsymbol{w_{3}}\colon$ $m_{4}^{*}$ ${m_{3}^{\dagger}}$ $m_{1}$ $\underline{m_{2}}$
@@ 2.2 Dataset, Models, and Setup
Preference Instances. We synthetically sample a set of $300$ preference profiles, partitioned into three sets of $100$ instances for each difficulty level, namely Easy ($n=10$ agents on each side of the market), Medium ($n=20$), and Hard ($n=50$). The preference profiles are sampled from two types of distributions Impartial Culture (IC) and Master-list (ML), each constituting $50$ questions at each difficulty level. An impartial culture (IC) is a well-studied probabilistic model for generating preference profiles in which every agent’s strict preference ranking is drawn independently and uniformly at random [4, 22]. It has been extensively studied in the context of economics, matching, and voting theory [7, 5, 68, 9]. A profile with a master-list (ML) is a highly structured preference profile in which all agents on one side of the market share exactly the same strict ranking over the agents on the other side. They represent the homogeneity in settings ranging from the labor market to organ allocation in healthcare [37, 6, 42] While an arbitrary instance generated by IC may admit exponentially many stable solutions [44], with a master-list, only a single unique stable solution exists, indicating a difficulty level proportional to the size of the space of stable outcomes. In Appendix C, we discuss a simpler version of the DA algorithm for computing stable solutions with ML instances.
@@
Matching Dataset. We curate a dataset comprising 2850 questions derived from the instances described above. [^2] Our novel dataset, along with code and preference instances, will be made publicly available. These questions cover four task categories, each applied to the same pool of profiles to ensure consistency: (i) generating stable solutions, given a preference profile ($300$ questions); (ii) instability resolution, given a profile and an unstable matching; (iii) instability detection, given a profile and a solution ($1050$ questions); and (iv) preference reasoning, given a single preference list or a profile ($900$ questions).
@@
Models. We select a representative suite of both open-source and closed-source models for evaluation. Since our benchmark is based on a reasoning task, we categorize models by their reasoning ability. We evaluate two basic models (those not specifically trained for reasoning), namely Llama-3.3-70B [21] and Gemini-2.0-Flash [59], and five reasoning models, namely Qwen-QwQ-32B [65], DeepSeek-70B (Llama-distilled) [17], OpenAI o3-mini [58], DeepSeek-R1 [17], and Gemini-2.5-Pro [16]. Among reasoning models, we classify the last three as advanced reasoning models, based on their SOTA performance on reasoning benchmarks [48].
@@
Prompting. The prompt for each task consists of the preference profile for a given instance, followed by task-specific instructions (e.g., computing the “proposer-optimal” matching, or resolving a given unstable matching). While we adopt the stable-marriage setting, considered by Gale and Shapley [27], where men propose to women, we show (in Appendix D) that the results do not change if a different setting—where workers are assigned tasks—is used. To scale up the verification of solution correctness, we instruct LLMs to adhere to a predefined answer format. Additionally, we allow LLMs two re-tries to correct solutions that are either invalid, partial, or do not adhere to the specified format. See Appendix H for details about the inference setup, and Appendix I for sample prompts.
@@ 2.3 Evaluation Criteria
Optimality/Pessimality Rate: This rate assesses the overlap between the model’s matching and a reference stable matching, thereby capturing how closely the model’s output mirrors the stepwise proposals and acceptances of a canonical algorithm. Formally, given two perfect matchings, $\mu$ and $\mu^{\prime}$, in a one-to-one market where each matching is viewed as a set of unordered pairs between agents, the Jaccard similarity is defined as $\text{JS}(\mu,\mu^{\prime})=\frac{|\mu\cap\mu^{\prime}|}{|\mu\cup\mu^{\prime}|}$. Then, we define optimality rate (OR) of a stable matching as its similarity to the proposer-optimal stable solution, which is unique.

### v2

@@ Abstract
The rise of Large Language Models (LLMs) has driven progress in reasoning tasks, from program synthesis to scientific hypothesis generation, yet their ability to handle ranked preferences and structured algorithms in combinatorial domains remains underexplored. We study matching markets, a core framework behind applications like resource allocation and ride-sharing, which require reconciling individual ranked preferences to ensure stable outcomes. We evaluate seven state‐of‐the‐art models on a hierarchy of preference‐based reasoning tasks—ranging from stable‐matching generation to instability detection, instability resolution, and fine-grained preference queries—to systematically expose their logical and algorithmic limitations in handling ranked inputs. Surprisingly, even top-performing models with advanced reasoning struggle to resolve instability in large markets, often failing to identify blocking pairs or execute algorithms iteratively. We further show that parameter-efficient fine-tuning (LoRA) significantly improves performance in small markets, but fails to bring about a similar improvement in large instances, suggesting the need for more sophisticated strategies to improve LLMs’ reasoning with larger-context inputs.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has positioned them as integral components in a wide range of reasoning-intensive tasks such as program synthesis, logical inference, mathematical problem solving, and scientific hypothesis generation, highlighting the importance of structured problem-solving capabilities. Despite their recent success in symbolic and logical reasoning, their capacity to reason over ranked preferences and to execute structured algorithms within combinatorial domains remains largely unexplored. Preference reasoning constitutes a foundational component in numerous domains, including economic contexts—e.g., auctions, voting systems, and market design—and in the architecture of pre-trained generative models using Reinforcement Learning from Human Feedback (RLHF) to capture and internalize human value judgments. These methods often have to execute algorithms on a large number of preference lists (either pairwise, partial, or complete rankings) to aggregate the rankings through constitutional AI [bai2022constitutional] or social choice theory [Conitzer+24].
Despite substantial progress, reasoning over preferences remains a non-trivial endeavor: ensuring transitivity [xu2025investigating, song2025benchmarking], accurately augmenting ordinal rankings [fish2023generative], and achieving coherent value alignment pose significant challenges. Without robust mechanisms for preference elicitation and the capacity to execute the requisite combinatorial procedures, even state-of-the-art LLMs may produce outputs that diverge from true human preferences [hosseini2025distributive] or fail to satisfy desirable properties [Fish2025EconEvals].
We consider matching markets, a domain that constitutes a fundamental class of problems underlying diverse applications—from healthcare resource allocation to ride-sharing platforms and recommender systems—and demand accurate comprehension of individual preferences and reconciliation of conflicting choices to guarantee system-wide stability. Matching markets are an ideal testbed for studying reasoning in AI models for two key reasons: First, they provide a structured platform for evaluating reasoning over ranked preferences and algorithmic thinking. They provide a framework that is axiomatically rich yet computationally tractable where solution quality (e.g., stability and efficiency) can be rigorously evaluated. Second, LLMs are increasingly utilized as black-box systems in a variety of economic, social, or medical settings to inform automated screening in recruitment pipelines [gan2024application], investigating market behavior [jia2024experimental], market clearing in ride-hailing platforms [junque2025simulating], and in general simulating economic interactions [horton2023large]. This makes it imperative to evaluate their ability in terms of computing “desirable” solutions from stakeholders’ preferences. Additionally, their potential to serve as interfaces between stakeholders and established black-box systems (e.g., the National Residents Matching Program [NRMP2024]) requires benchmarking their ability to reason about provided solutions and address aspects that users might find undesirable.
@@ 1.1 Our Results
Figure 1: The framework for reasoning with ranked preferences through matching markets.
We focus on four preference-based tasks: (i) generating stable solutions, requiring LLMs to produce stable matchings directly from ranked inputs, (ii) instability resolution, demanding preference reasoning to transform unstable matchings to stable ones, (iii) instability detection, in which models detect blocking pairs within a proposed solution, and (iv) preference reasoning, assessing nuanced query answering over ranked lists. We evaluate seven large language models with varying reasoning capabilities, including basic models, those with some limited reasoning, and advanced reasoning models. Our methodology and results are summarized in Figure˜1.
Benchmark. We introduce a benchmark with instances and questions aimed at evaluating the above tasks involving reasoning over ranked preferences. These tasks are categorized into three levels of difficulty—Easy, Medium, and Hard—based on problem size. Each task utilizes ranked preferences sampled from two statistical distribution models: Impartial Culture (IC) and Master-list (ML).
@@
Instability Detection and Resolution. We find that LLMs frequently make mistakes in determining whether solutions are stable, with hallucinations about blocking pairs being the most common among basic models. Additionally, LLMs’ ability to correct unstable solutions is (at best) as good as their ability to generate them from scratch, in some cases making the provided incorrect solutions worse.
@@ 1.2 Related Work
Reasoning Capabilities of LLMs. Mathematical problem solving has been a key area of focus in evaluating the reasoning ability of LLMs, through a variety of benchmarks such as [HendrycksMATH2021, HendrycksMMLU2021, CobbeGSM8K2021, Mirzadeh2024GSMSymbolic]. LLMs have also demonstrated remarkable capabilities on coding benchmarks such as SWE-Bench [Jiminez2024SWEBench] and CodeForces. As SOTA benchmark scores improve, recent work studies whether these improvements reflect genuine logical reasoning through benchmarks assessing logical consistency [Liu2024Logic] and rule understanding/execution/planning [Gui2024Rule]. Furthermore, the recent rise of LLM agents has increased interest in benchmarking LLMs’ causal reasoning [Haoang2024Causal] and strategic planning abilities [Duan2024GTBench, Tang2025DSGBench, Jia2024Strategic]. Additionally, the emergence of reasoning models has led to benchmarks evaluating these models’ improved reasoning and planning abilities [ChenReasoningEra2025, LiSystem22025].
Enhancing Reasoning Capabilities of LLMs. Specialized prompting strategies like Chain-of-Thought (CoT) [wei2022chain], Tree-of-Thought (ToT) [Yao2023ToT], and Graph-of-Thought (GoT) [Besta2024GoT] have performance abilities on a variety of reasoning benchmarks. Fine-tuning has also been demonstrated to improve CoT in model outputs [Zhang2024Chain], as well as economic rationality [chen2023emergence] and abstract reasoning [Globersons2024Learning]. Additionally, instruction-tuning has been shown to enhance reasoning in several works [Luo2024Improve, Vaillancourt_2024, Cai2024System2, Li2024MuggleMath]. More advanced techniques build upon CoT [Tian2024Improvement, Zhou2024Discover], or utilize multi-agent architectures that leverage cooperative LLMs [Zhou2024Reso, Trencsenyi2025Reasoning]. More recently, reinforcement-learning (GRPO) has been used to improve model reasoning, the most popular example being the Deepseek-R1 reasoning model [DSR1].
LLMs in Social and Economic Decision Making. While still an emerging area of research, multiple works have focused on the collective decision-making capabilities of LLMs. One particular area of interest is the use of LLMs in preference elicitation [Soumalis2025Elicitation, Huang2025Accelerated]. Fish2025EconEvals benchmark the ability of models to learn and strategize in unknown economic environments using deliberate exploration. Another notable avenue of work is the study of how well LLMs can represent humans in collective decision-making, an understudied component of LLM alignment [Yang2024Voting, hosseini2025distributive].
@@ 2.1 Problem Formulation
A two-sided matching market consists of two disjoint sets of agents (e.g., riders and drivers, freelancers and job requesters, and content creators and ads) denoted by $M$ and $W$, where $|M|=|W|=n$. The preference list of an agent $i$, denoted by $\succ_{i}$, is a ranked order list over the agents on the other side. A preference profile, $\succ$, denotes the collection of preferences of all agents. We write $w_{1}\succ_{m}w_{2}$ and $m_{1}\succ_{w}m_{2}$ to denote that $m$ prefers $w_{1}$ to $w_{2}$ and $w$ prefers $m_{1}$ to $m_{2}$ respectively. In this paper, we primarily consider the standard model, which assumes a complete and strict preference list (no ties) and aims at finding a one-to-one matching between the agents in two sets. [^1] This is the standard model considered by the seminal works of gale1962college and knuth1997stable.
Matching and Stability. A matching is a function $\mu:M\cup W\rightarrow M\cup W$ such that $\mu(m)\in W$ for all $m\in M$, $\mu(w)\in M$ for all $w\in W$, and $\mu(m)=w$ if and only if $\mu(w)=m$. Given a matching $\mu$, a blocking pair with respect to the preference profile $\succ$ is a pair $(m,w)$ who prefer each other over their assigned partners in $\mu$, i.e., $w\succ_{m}\mu(m)$ and $m\succ_{w}\mu(w)$. A matching is said to be stable if it does not have any blocking pairs. Given an instance of the problem, the set of all possible stable solutions forms a distributive lattice and can be exponential in size [knuth1997stable].
In their seminal work, gale1962college proposed an iterative procedure—the deferred acceptance algorithm (DA)—that always guarantees to find a stable solution. It proceeds by a series of proposals and rejections. In the initial proposal phase, each of the unmatched agents on one side (aka proposers) proposes to their favorite agent from the other side (aka receivers) according to their preference list. In the subsequent rejection phase, each agent on the receiving side tentatively accepts its preferred proposal, rejecting the others. The algorithm terminates when no further proposals can be made. The details of this algorithm can be found in Appendix˜C. The underlined solution in Example˜1 is simultaneously optimal for the proposing side and pessimal for the receiving side [mcvitie1971stable]. We refer to the former as the Optimal matching and the latter as the Pessimal matching.
@@
[…]
@@ 2.2 Dataset, Models, and Setup
Preference Instances. We synthetically sample a set of $300$ preference profiles, partitioned into three sets of $100$ instances for each difficulty level, namely Easy ($n=10$ agents on each side of the market), Medium ($n=20$), and Hard ($n=50$). The preference profiles are sampled from two types of distributions Impartial Culture (IC) and Master-list (ML), each constituting $50$ questions at each difficulty level. An impartial culture (IC) is a well-studied probabilistic model for generating preference profiles in which every agent’s strict preference ranking is drawn independently and uniformly at random [black1958theory, eugeciouglu2013impartial]. It has been extensively studied in the context of economics, matching, and voting theory [brilliantova2022fair, boehmer2024guide, TsetlinIC, Caragiannis2024IC]. A profile with a master‐list (ML) is a highly structured preference profile in which all agents on one side of the market share exactly the same strict ranking over the agents on the other side. They represent the homogeneity in settings ranging from the labor market to organ allocation in healthcare [irving2008stable, Bredereck2021MasterList, NaoyukiMLMatroid] While an arbitrary instance generated by IC may admit exponentially many stable solutions [knuth1997stable], with a Master-list, only a single unique stable solution exists, indicating a difficulty level proportional to the size of the space of stable outcomes. In Appendix˜C, we discuss a simpler version of the DA algorithm for computing stable solutions with ML instances.
@@
Matching Dataset. We curate a dataset comprising 2850 questions derived from the instances described above. These questions cover four task categories, each applied to the same pool of profiles to ensure consistency: (i) generating stable solutions, given a preference profile ($300$ questions); (ii) instability resolution, given a profile and an unstable matching; (iii) instability detection, given a profile and a solution ($1050$ questions); and (iv) preference reasoning, given a single preference list or a profile ($900$ questions).
@@
Models. We select a representative suite of both open-source and closed-source models for evaluation. Since our benchmark is based on a reasoning task, we categorize models by their reasoning ability. We evaluate two basic models (those not specifically trained for reasoning), namely Llama-3.3-70B [Dubey2023Llama] and Gemini-2.0-Flash [sundar_pichai_2024], and five reasoning models, namely Qwen-QwQ-32B [qwen_paper], DeepSeek-70B (Llama-distilled) [DSR1], OpenAI o3-mini [openai_o3mini_2025], DeepSeek-R1 [DSR1], and Gemini-2.5-Pro [deepmind_2025]. Among reasoning models, we classify the last three as advanced reasoning models, based on their SOTA performance on reasoning benchmarks [LiSystem22025].
@@
Prompting. The prompt for each task consists of the preference profile for a given instance, followed by task-specific instructions (e.g., computing the “proposer-optimal” matching, or resolving a given unstable matching). While we adopt the stable-marriage setting, considered by gale1962college, where men propose to women, we show (in Appendix˜D) that the results do not change if a different setting—where workers are assigned tasks—is used. To scale up the verification of solution correctness, we instruct LLMs to adhere to a predefined answer format. Additionally, we allow LLMs two re-tries to correct solutions that are either invalid, partial, or do not adhere to the specified format. See Appendix˜I for details about the inference setup, and Appendix˜J for sample prompts.
@@ 2.3 Evaluation Criteria
Optimality/Pessimality Rate: This rate assesses the overlap between the model’s matching and a reference stable matching, thereby capturing how closely the model’s output mirrors the stepwise proposals and acceptances of a canonical algorithm. Formally, given two perfect matchings, $\mu$ and $\mu^{\prime}$, in a one‐to‐one market where each matching is viewed as a set of unordered pairs between agents, the Jaccard similarity is defined as $\text{JS}(\mu,\mu^{\prime})=\frac{|\mu\cap\mu^{\prime}|}{|\mu\cup\mu^{\prime}|}$. Then, we define optimality rate (OR) of a stable matching as its similarity to the proposer-optimal stable solution, which is unique.